OpenStudy (anonymous):

angle 1 and 2 are called angles

OpenStudy (alexandervonhumboldt2):

yeah all angles are angles

OpenStudy (anonymous):

not done lol

OpenStudy (anonymous):



OpenStudy (pawanyadav):

Angle 1 and angle 2 forms a linear pair

OpenStudy (pawanyadav):

Means 1+2=180° So what are they?

OpenStudy (alexandervonhumboldt2):



OpenStudy (mathmale):

Yes, and a linear pair add up to 180 degrees.

OpenStudy (anonymous):

complementary

OpenStudy (pawanyadav):

As their sum is 180 you can also say them supplementary.

OpenStudy (anonymous):

forreal

OpenStudy (mathmale):

Can you be more specific in indicating that 1 and 2 represent angles? Means 1+2=180° is pretty inaccurate; 1+2 = 3, right? What symbolism could you use to show that 1 and 2 represent angles #1 and #2?

OpenStudy (pawanyadav):

No, sum of complementary=90°

OpenStudy (mathmale):

So "complementary" does not apply in this particular problem.

OpenStudy (mathmale):

Hey, Cam, what are you looking for? What are the instructions for this problem?

OpenStudy (anonymous):

to find if it is a vertical right complentary or supplementary angle

OpenStudy (mathmale):

Sorry to be picky, but I need to know to what "it" refers to. Angles (plural) can be supplementary; single angles cannot.

OpenStudy (anonymous):

it doesnt tell

OpenStudy (anonymous):

single angle it doesnt tell just says 1 and 2 is a type of angle and i showed you the picture

OpenStudy (anonymous):

but mathmale is right